# Clustering in Networks with Time-varying Nodal Attributes

### v1

@@ 4.3 Word Co-occurrence Network
We also analyze word usage from the novel David Copperfield by Charles Dickens in 1850. newman2006finding constructed a word co-occurrence network containing $112$ common adjectives and nouns from the book. The network nodes correspond to different words, and an undirected edge between two words indicates whether they occur adjacent to one another at any sentence within the novel.
@@
$Y_{t,i}=\frac{C_{t,i}-\bar{C}_{i}}{\hat{\sigma}_{i}},\hskip 18.49988ptt=1,\dots,n,\qquad i\in\mathcal{V}_{\text{word}}.$
where $\bar{C}_{i}$ and $\hat{\sigma}_{i}$ denote the respective sample mean and standard deviation of the $i$th’s word series. This standardization ensures that the analysis focuses on the relative temporal dynamics of word usage rather than count magnitudes.
Figure 6(a) displays our clustering. Node shapes indicate the true linguistic labels (noun or adjective), while colors depict estimated clusters. The results demonstrate that most nouns (⚫) fall in a cluster (Cluster 1) and most adjectives (◼) into another (Cluster 2). The other two clusters are small, each containing no more than three words (and connected only to a single node in the remaining network). In particular, Cluster 3 contains the two words “aunt” and “family”, reflecting the prominence of family as a central theme in the novel and the pivotal role of Aunt Betsey. Unlike David’s mother, whose influence fades with her early passing, Aunt Betsey becomes the loving guardian, shaping David’s values and providing unwavering support throughout his growth.
(a) GFL
(b) $k$-means ($k=2$)
Figure 6: Clusters from GFL and $k$-means methods. Node shapes indicate whether a word is an adjective or noun, while colors depict cluster members.
Figure 6(b) displays the $k$-means clustering. Based on the linguistic nature (noun or adjective) of the words, two clusters were chosen. The $k$-means method fails to discover coherent clusters, likely because it disregards the co-occurrence structure delineated by the graph. As a result, nouns and adjectives intermix in the clusters. This contrast underscores the advantage of our method.
@@ 5 Discussion
This manuscript introduced a decoder-only latent space framework for clustering nodes in networks that have an observed time series at each node. By combining node-specific Gaussian priors, graph-fused LASSO regularization, and a neural network decoder, series and network dynamics are integrated into low-dimensional representations that facilitate clustering. Simulations and applications to California county temperatures and word co-occurrence network produced good clustering performance.
Avenues for future development are apparent. Foremost, the type of nodal data could change. Multiple series, functional data, or even a simple multivariate attribute are feasible. Our likelihood would need modification to accommodate such scenarios. The assumption of independent attributes across differing nodes could also be relaxed. On the network front, nodal relations often have a degree of strength, perhaps a continuous value in lieu of zero-one connections (krivitsky2012exponential; wilson2017stochastic). Adapting our LASSO regularization to handle such structure would further improve clustering. Time-varying network edges can also arise (malinovskaya2023statistical). Methods to accommodate both nodal and dyadic attributes over time would extend applicability of our methods.
6 Data and Code Availability
7 Disclosure Statement
The authors report that there are no competing interests to declare.
SUPPLEMENTARY MATERIAL
@@ A.1 Updating $\bm{\mu}$ and $\bm{\phi}$
Note that the expectation in the gradient is now with respect to the posterior density function $P(\bm{Z}_{i}|\bm{Y}_{i})\propto P(\bm{Y}_{i}|\bm{Z}_{i})\times P(\bm{Z}_{i})$. Furthermore, denote the objective function in Equation (4) as
@@ A.2 Updating $\bm{\nu}$
$=\left(\frac{\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}-\frac{\lambda}{\gamma}}{\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}}\right)^{-1}\big[\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\big]$
@@
we obtain the condition that $\bm{\nu}_{i,j}$ becomes zeros when $\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}\leq\lambda$. Therefore, we can iteratively update $\bm{\nu}_{i,j}$ for $(i,j)\in\mathcal{E}$ with the following equation:
@@
A.3 Computation Complexity
The algorithm to solve (2) via ADMM is presented in Algorithm 1. The computation complexity is at least of order $O\big(A(|\mathcal{V}|skd+B|\mathcal{V}|sn+|\mathcal{V}|sd+D|\mathcal{E}|d)\big)$; additional gradient computation time is required for neural networks in sub-routines.
Elaborating, in each of $A$ iterations of ADMM, Langevin dynamics generate $s$ samples for each $i$, each requiring $k$ MCMC steps (and has dimension of $d$). Next, the decoder parameters $\bm{\phi}$ are updated via $B$ iterations of the Adam optimizer across all nodes using the $s$ generated samples. The output of neural networks has length $n$ for each node. Then, the node-specific prior parameters $\bm{\mu}_{i}$ are updated in closed form at each iteration using the MCMC samples, which incurs a computational cost that is linear in $|\mathcal{V}|sd$. Finally, the slack variables $\bm{\nu}_{i,j}$ are updated by block coordinate descent, repeated for $D$ iterations, resulting in a complexity of $O(D|\mathcal{E}|d)$. The scaled dual variables $\bm{w}_{i,j}$ are updated in closed form and effectively do not increase computational complexity.
@@ A.3 Computation Complexity
10:   Set $\widetilde{\bm{\nu}}^{(1)}=\bm{\nu}^{(a)}$
11:   for $d=1,\dots,D$ do
12:    Let $\widetilde{\bm{\nu}}_{i,j}^{(d+1)}$ be updated according to (9) for $(i,j)\in\mathcal{E}$
13:   end for
14:   Set $\bm{\nu}^{(a+1)}=\widetilde{\bm{\nu}}^{(d+1)}$
15:   $\bm{w}_{i,j}^{(a+1)}=\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}^{(a+1)}+\bm{w}_{i,j}^{(a)},\ \forall\ (i,j)\in\mathcal{E}$
16:end for
@@
18:Output: learned prior parameters $\hat{\bm{\mu}}$
A.4 Model Selection
The optimization problem in (2) involves the tuning parameter $\lambda$. Given a list of candidate $\lambda$ values, model selection is performed via cross validation. Specifically, the observed data is split into training and testing sets via nodes. Let $\mathcal{V}_{\text{test}}$ be the node set of the testing data. During training, we replace the node $i$ data $\bm{Y}_{i}$ with the element-wise average $\bar{\bm{Y}}_{i}$ from all neighbors. The node $i$ data is held out for testing when $i\in\mathcal{V}_{\text{test}}$.
$\sum_{i\in\mathcal{V}_{\text{test}}}\ln\big(P(\bm{Y}_{i})\big)=\sum_{i\in\mathcal{V}_{\text{test}}}\ln\Big[\int P(\bm{Y}_{i}|\bm{Z}_{i})P(\bm{Z}_{i})d\bm{Z}_{i}\Big]\approx\sum_{i\in\mathcal{V}_{\text{test}}}\ln\left[\frac{1}{s}\sum_{u=1}^{s}\big[P(\bm{Y}_{i}|\bm{Z}_{u,i})\big]\right]$
@@ Appendix B Additional Details of Experiments
For particulars, $k$-means constructs clusters by minimizing within-cluster squared distances. DTW method measures similarities between time series by aligning them via a non-linear time warping. CASC method leverages the nodal covariate $\bm{X}$ via $\bm{L}_{\tau}\bm{L}_{\tau}+\alpha\bm{X}\bm{X}^{\top}$ to improve spectral clustering performance; here, $\bm{L}_{\tau}$ is the regularized graph Laplacian. SDP method converts clustering issues into a convex semi-definite procedure, enabling efficient approximation of the optimal partition. Specifically, SDP maximizes the inner product between $\bm{A}+\lambda\bm{K}$ and a cluster label matrix; here, $\bm{A}$ is an adjacency matrix and $\bm{K}\in\mathbb{R}_{+}^{N\times N}$ is a kernel matrix whose $(i,j)$th entry quantifies the similarity between covariates from nodes $i$ and $j$. NAC method incorporates node covariates into spectral clustering through the network-adjusted covariate matrix $\bm{A}\bm{X}+\bm{D}_{\alpha}\bm{X}\in\mathbb{R}^{N\times p}$ with diagonal weight matrix $\bm{D}_{\alpha}$ and covariate matrix $\bm{X}$. The SCORE method identifies communities through the ratios between the first leading eigenvector and the other leading eigenvectors.
For competitors, the true number of clusters from the simulated data and their default configurations are used during implementation. For GFL, we run our experiments with A100 GPU. Table 4 summarizes the number of model parameters and data points, when the node size is $N=120$, the latent dimension is $d=3$, and the sequence length is $n=100$. The neural network decoder consists of two hidden layers, each containing $32$ neurons.
Table 4: Number of parameters and data points.
Notation Description Quantity
$\bm{\phi}\$Neural network parameters $32\times 3+32\times 32+100\times 32\ +$ $32\times 1+32\times 1+100\times 1=4{,}484$
$\{\bm{\mu}_{i}\}_{i\in\mathcal{V}}$ Prior means $120\times 3=360$
$\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ Nodal series $120\times 100=12{,}000$
B.1 Visualization
@@
Figure 7: Illustration of estimated $\hat{\bm{\mu}}\in\mathbb{R}^{N\times d}$ for a realization from Scenarios 1 - 3, where the latent dimension is $d=3$. Each row, from top to bottom, corresponds to a scenario; each column, from left to right, displays a pairwise projection of the three latent dimensions. The node colors depict the ground truth labels.
@@
B.2 Sensitivity Analysis
@@
Table 5: Sample means (one standard deviation) of our evaluation metrics for varied latent dimension $d\in\{3,5,7,10\}$ and Scenarios (Scen.) 1-3. The best metric is in bold.
Scen., NN $d$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ HOM $\uparrow$ COM $\uparrow$ PUR $\uparrow$

### v2

@@ 5.3 Word Co-occurrence Network
We also analyze word usage from the novel David Copperfield, written by Charles Dickens in 1850. (Newman, 2006) constructed a word co-occurrence network containing $112$ common adjectives and nouns from the book. The network nodes correspond to different words, and an undirected edge between two words indicates whether they occur adjacent to one another at a sentence within the novel.
@@
$Y_{t,i}=\frac{C_{t,i}-\bar{C}_{i}}{\hat{\sigma}_{i}},\qquad t=1,\dots,n,\quad i\in\mathcal{V}_{\text{word}},$
where $\bar{C}_{i}$ and $\hat{\sigma}_{i}$ denote the respective sample mean and standard deviation of the word’s series. This standardization ensures that the analysis focuses on the relative temporal dynamics of word usage rather than count magnitudes.
Figure 6 displays our clustering. Node shapes indicate the true linguistic labels (noun or adjective), while colors depict estimated clusters. The results demonstrate that most nouns (⚫) fall in a cluster (Cluster 1) and most adjectives (◼) into another (Cluster 2). The other two clusters are small, each containing no more than three words and connected only to a single node in the remaining network. Similar to (Newman and Leicht, 2007), our method recovers the bipartite structure without being explicitly prompted. In particular, Cluster 3 contains two words “aunt” and “family”, reflecting the prominence of family as a central theme in the novel and the pivotal role of Aunt Betsey. Unlike David’s mother, whose influence fades with her early passing, Aunt Betsey becomes the loving guardian, shaping David’s values and providing unwavering support throughout his growth.
Figure 6: Word clustering from GFL. Node shapes indicate whether a word is an adjective or noun, while colors depict cluster members.
@@ 6 Discussion
This manuscript presents a decoder-only latent space model for clustering nodes in networks with node-level observed time series. By combining node-specific Gaussian priors, graph-fused LASSO regularization, and a neural network decoder, graph structures and attribute dynamics are integrated into low-dimensional representations that facilitate clustering. Simulations and applications to California county temperatures and word co-occurrence network produce good clustering performance.
Avenues for future development are apparent. Foremost, the type of nodal data could change. Multiple series, functional data, or multivariate attributes are feasible. The likelihood would need modification to accommodate such scenarios. The assumption of independent attributes across nodes could also be relaxed. On the network front, dyadic relations often have a degree of strength, perhaps continuous values in lieu of zero-one connections (Krivitsky, 2012; Wilson et al., 2017). Adapting our LASSO regularization to handle such weights would further improve clustering. Time-varying network edges can also arise (Malinovskaya et al., 2023). Methods to accommodate both nodal and dyadic attributes over time would extend applicability of our methods.
@@ A.1 Updating $\bm{\mu}$ and $\bm{\phi}$
Note that the expectation in the gradient is now with respect to the posterior density function $P(\bm{Z}_{i}|\bm{Y}_{i})\propto P(\bm{Y}_{i}|\bm{Z}_{i})\times P(\bm{Z}_{i})$. Furthermore, denote the objective function in (4) as
@@ A.2 Updating $\bm{\nu}$
$=\left(\frac{\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}-\frac{\lambda}{\gamma}}{\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}}\right)\big[\gamma(\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j})\big]$
@@
we obtain the condition that $\bm{\nu}_{i,j}$ becomes zeros when $\gamma\|\bm{\mu}_{i}-\bm{\mu}_{j}+\bm{w}_{i,j}\|_{2}\leq\lambda$. Therefore, we can update $\bm{\nu}_{i,j}$ for $(i,j)\in\mathcal{E}$ with the following equation:
@@
A.3 Neural Decoder
For flexibility of decoder $P_{\bm{\phi}}(\bm{Y}_{i}|\bm{Z}_{i})$, the mean $\bm{h}_{\bm{\phi}}(\bm{\cdot})$ is parameterized by a three-layer ReLU neural network. Specifically, $\bm{h}_{\bm{\phi}}(\bm{\cdot})$ is taken as a three-layer rectified linear unit (ReLU) neural network:
where $\bm{W}_{1}\in\mathbb{R}^{h_{1}\times d},\bm{W}_{2}\in\mathbb{R}^{h_{2}\times h_{1}}$, and $\bm{W}_{3}\in\mathbb{R}^{n\times h_{2}}$ are weights and $\bm{b}_{1}\in\mathbb{R}^{h_{1}\times 1},\bm{b}_{2}\in\mathbb{R}^{h_{2}\times 1}$, and $\bm{b}_{3}\in\mathbb{R}^{n\times 1}$ are bias parameters. Here, $\mathrm{ReLU}(x)=\max(0,x)$ is the element-wise ReLU activation function at $x$, and model parameters are collected into $\bm{\phi}$.
Table 4 summarizes the number of model parameters and data points, when the node size is $|\mathcal{V}|=120$, the latent dimension is $d=3$, and the sequence length is $n=100$. The neural network decoder consists of two hidden layers, each containing $32$ neurons. The three-layer ReLU neural networks can be expressive without being over-parameterized.
Table 4: Number of parameters and data points.
Notation Description Quantity
$\bm{\phi}\$Neural networks $32\times 3+32\times 32+100\times 32\ +$ $32\times 1+32\times 1+100\times 1=4{,}484$
$\{\bm{\mu}_{i}\}_{i\in\mathcal{V}}$ Prior means $|\mathcal{V}|\times 3=360$
$\{\bm{Y}_{i}\}_{i\in\mathcal{V}}$ Nodal series $|\mathcal{V}|\times 100=12{,}000$
The negative log-likelihood that involves latent variable $\bm{Z}_{i}\in\mathbb{R}^{d}$ and observed series $\bm{Y}_{i}\in\mathbb{R}^{n}$ is lower bounded by
$-\ln\big(P(\bm{Y}_{i})\big)$ $=-\ln\left[\int\frac{1}{(2\pi)^{n/2}}\exp\left(-\frac{1}{2}\|\bm{Y}_{i}-\bm{h}_{\bm{\phi}}(\bm{Z}_{i})\|_{2}^{2}\right)\cdot\frac{1}{(2\pi)^{d/2}}\exp\left(-\frac{1}{2}\|\bm{Z}_{i}-\bm{\mu}_{i}\|_{2}^{2}\right)d\bm{Z}_{i}\right]$
$\geq-\ln\left[\frac{1}{(2\pi)^{(n+d)/2}}\int\exp\left(-\frac{1}{2}\|\bm{Z}_{i}-\bm{\mu}_{i}\|_{2}^{2}\right)d\bm{Z}_{i}\right]$ (12)
$\geq-\ln\left[\frac{1}{(2\pi)^{n/2}}\right]$
$\geq\frac{n}{2}\ln(2\pi)$
where (12) follows because $\exp(-\frac{1}{2}x^{2})\leq 1$. While the likelihood function of some latent space models can be unbounded (Mattei and Frellsen, 2018), the minimization problem in (2) is valid and meaningful.
A.4 Computation Complexity
The algorithm to solve (3) via ADMM is presented in Algorithm 1. The computation complexity is at least of order $O(A(|\mathcal{V}|skd+B|\mathcal{V}|sn+|\mathcal{V}|sd+|\mathcal{E}|d))$; additional gradient computation for neural networks is required in sub-routines.
Elaborating, in each of $A$ iterations of ADMM, Langevin dynamics generate $s$ samples for each $i$, each requiring $k$ MCMC steps (and has dimension of $d$). Next, the decoder parameters $\bm{\phi}$ are updated via $B$ iterations of the Adam optimizer across all nodes using the $s$ generated samples. The output of neural networks has length $n$ for each node. Then, the node-specific prior parameters $\bm{\mu}_{i}$ are updated in closed form at each iteration using the MCMC samples, which incurs a computational cost that is linear in $|\mathcal{V}|sd$. Finally, the slack variables $\bm{\nu}_{i,j}$ are updated, resulting in a complexity of $O(|\mathcal{E}|d)$. The scaled dual variables $\bm{w}_{i,j}$ are updated in closed form with similar computation complexity.
@@ A.4 Computation Complexity
10:  Let $\bm{\nu}_{i,j}^{(d+1)}$ be updated according to (8) for $(i,j)\in\mathcal{E}$
11:  $\bm{w}_{i,j}^{(a+1)}=\bm{\mu}_{i}^{(a+1)}-\bm{\mu}_{j}^{(a+1)}-\bm{\nu}_{i,j}^{(a+1)}+\bm{w}_{i,j}^{(a)},\ \forall\ (i,j)\in\mathcal{E}$
12:end for
@@
14:Output: learned prior parameters $\hat{\bm{\mu}}$
@@ Appendix B Additional Details of Experiments
B.1 Implementation Details
To simulate realistic network structures, block and grid graphs are considered. For the block graphs in Scenarios 1 and 3, the edge density is roughly $2|\mathcal{E}||\mathcal{V}|^{-2}\approx 0.4$. In Scenario 2, the edge density for a $12\times 12$ grid is about $2|\mathcal{E}||\mathcal{V}|^{-2}\approx 0.051$, and the edge density for a $14\times 14$ grid is about $2|\mathcal{E}||\mathcal{V}|^{-2}\approx 0.038$. In reality, many networks are considered small and sparse, and the objective of researchers is to draw insights from limited information.
For particulars of competitor methods, $k$-means constructs clusters by minimizing within-cluster squared distances. DTW method measures similarities between time series by aligning them via a non-linear time warping. CASC method leverages the nodal covariate $\bm{X}$ via $\bm{L}_{\tau}\bm{L}_{\tau}+\alpha\bm{X}\bm{X}^{\top}$ to improve spectral clustering performance; here, $\bm{L}_{\tau}$ is the regularized graph Laplacian. SDP method converts clustering issues into a convex semi-definite procedure, enabling efficient approximation of the optimal partition. Specifically, SDP maximizes the inner product between $\bm{A}+\lambda\bm{K}$ and a cluster label matrix; here, $\bm{A}$ is an adjacency matrix and $\bm{K}\in\mathbb{R}_{+}^{N\times N}$ is a kernel matrix whose $(i,j)$th entry quantifies the similarity between covariates from nodes $i$ and $j$. NAC method incorporates node covariates into spectral clustering through the network-adjusted covariate matrix $\bm{A}\bm{X}+\bm{D}_{\alpha}\bm{X}\in\mathbb{R}^{N\times p}$ with diagonal weight matrix $\bm{D}_{\alpha}$ and covariate matrix $\bm{X}$. The SCORE method identifies communities through the ratios between the first leading eigenvector and the other leading eigenvectors.
For competitors, the true number of clusters from the simulated data and their default configurations are used during implementation. For GFL, we run our experiments with A100 GPU. As the objective function in (3) is highly non-convex due to the neural networks and sub-routines requiring MCMC sampling, specifying standard convergence criteria is impractical. The algorithm is implemented for a preset number of iterations, as an early stopping strategies to avoid over-fitting, which is commonly used in machine learning.
B.2 Visualization
@@
Figure 7: Illustration of estimated $\hat{\bm{\mu}}\in\mathbb{R}^{N\times d}$ for a realization from Scenarios 1 - 3, where the latent dimension is $d=3$. Each row, from top to bottom, corresponds to a scenario. Each column, from left to right, displays a pairwise projection of the three latent dimensions. The node colors depict the ground truth labels.
@@
B.3 Results from Competitors
Figure 10 displays the California county clusterings from the competitor methods. Ten clusters are selected to be comparable to our result; if the numbers of clusters for competitors are selected by silhouette scores, fewer clusters and unrealistic results arise. Overall, the clusterings are less geographically coherent. For the $k$-means, DTW, and SDP methods, Inyo County in the Owens Valley is grouped with the Desert South cluster. The DTW, CASC, and SCORE methods produce clusters that span both costal and inland regions. For the NAC method, San Diego County and Orange County are placed in the Desert South cluster, while Kern County in the Central Valley is grouped with the Lower Coastal region. For the $k$-means and CASC methods, the Northeast Mountain counties mix with the Coastal Northwest counties. With understanding of local geography, these results are less plausible and difficult to justify.
(a) $k$-means
(b) DTW
(c) CASC
(d) SDP
(e) NAC
(f) SCORE
Figure 10: California county clustering from competitor methods with ten clusters.
Figure 11 displays the clusterings of the word co-occurrence network from competitors. Based on the linguistic nature (noun or adjective) of the words, two clusters were chosen. Most competitors fail to discover coherent clusters, resulting in substantial mixing of nouns and adjectives. The $k$-means and DTW methods disregard the co-occurrence structure delineated by the graph, and the SCORE method overlooks the narrative dynamics across chapters. The SDP and NAC methods may integrate both sources of information improperly to exhibit degraded clustering performance. In contrast, our method yields more linguistically coherent partitions.
(a) $k$-means
(b) DTW
(c) CASC
(d) SDP
(e) NAC
(f) SCORE
Figure 11: Word clustering from competitor methods. Node shapes indicate whether a word is an adjective or noun, while colors depict cluster members.
B.4 Sensitivity Analysis
@@
Table 5: Sample means (one standard deviation) of our evaluation metrics for varied latent dimension $d\in\{3,5,7,10\}$ and Scenarios 1-3. The best metric is in bold.
Scenario, NN $d$ NMI $\uparrow$ ARI $\uparrow$ ACC $\uparrow$ HOM $\uparrow$ COM $\uparrow$ PUR $\uparrow$